Honework Week 5 Word t References Mailings Review View Gammarly Оте mewhat you want to do as .之. a- L. 1-1-1 Norrmal-I 1 NoSpec. 프 . . Heading 1 Heading 2 Title Subtitle Subtle Em.. Em phace Patacrach NEWS WIRE R&D INCENTIVES ES Two Drug Firms Agree to Settle Pricing Suit ALBANY, NY, Jan, 2 Z-Two drug companies have agreed to pay $80 million to settle allegations that they conspired to keep a cheaper, generic version of a blood pressure medication off the markest Under the settlement announced today Aventis Pharmaceuticals Inc. and Andrx Corp. will pay that amount to states, insurance companies, and consumers nationwide Consumers paid too much for the drugs Cardizem CD and its generic equivalents because the companies conspired y the marketing of cheaper competitors, said New York state Attorney General Eliot L. Spitzer Spitzer said that in 1998, the German pharmaceutical gant Hoechst-which merged with Rhone-Poulenc in 1999 to form Aventis-paid Andrx just under $100 milion to not market a generic form of Cardzem CD for 11 months. The agreement was to be renewed annually, he said n form of anticompetitive behavior kept the drug financially out of the reach of countless people Spitzer said Consumer groups have said that Cardizem sales total about $700 million a year domestically Users of C were paying about $73 a month for the drug when a generic cost ab Michael Gormley Source: Associated Press, January 28, 2003. Used with permission of The Associated Press. C 0 2003. All rights How much profit per year per user might the producers of Cardizem have been making if their average total costs were equal to $55 a month? per year per user References
Answer #1

In this case,

Payment per month by users of cardizem=P=$73

Average Cost cardizem=ATC=$55

Profit per user per month=Bm=P-ATC=73-55=$18

Profit per user per year=12*Bm=12*$18=$216